# SemNav: A Semantic Segmentation-Driven Approach to Visual Semantic Navigation

### v1

@@ III-B SemNav model
The architecture depicted in Figure 3 comprises several key components. To process the semantic segmentation input, we first transform it into an RGB format. Each pixel is assigned a semantic segmentation label, which is then mapped to a specific color in the RGB space using the semantic segmentation sensor integrated into the SemNav dataset. The encoded input is subsequently processed by a ResNet50 architecture [43], which extracts high-level visual features for downstream tasks. In our experiments, we evaluate two types of initialization for the ResNet50: (1) a random initialization and (2) an initialization using DINO [44]. The input corresponding to the target object category is converted into a one-hot vector and processed through an embedding layer. GPS and compass observations are directly passed through a fully connected layer. All these processed inputs are concatenated and subsequently fed into a Gated Recurrent Unit (GRU) network [45]. This recurrent network is responsible for predicting the discrete action distribution that forms the output of our model.
@@ IV-A Experimental settings and evaluation metrics
Our main goal in this work is to develop VSN solutions capable of navigating efficiently both in the virtual environments where they are trained and in the real world. For this purpose, we have designed two types of experiments.
@@
For this experimental evaluation, we addressed the ObjectNav task detailed in the Habitat 2023 challenge [16]. We used the official HM3D [21] training and validation sets. For our SemNav model we employed the two semantic segmentation sensors detailed in Section III-A. For evaluation metrics, we followed the standard proposals for the ObjectNav problem [16]. Specifically, these metrics include: 1) the Success Rate (SR), defined as the percentage of navigation episodes successfully completed by the model; 2) the Shortest Path Length (SPL), which compares the distance traveled by the agent during the episode with the shortest possible path length to the target object; 3) the average number of collisions the agents experience with the environment (C); and 4) the average distance to the goal (DTG), which is the mean distance of the agent from the nearest target object at the end of a navigation episode.
@@
For real-world experiments, the evaluations were conducted in a domestic environment, whose floor plan is shown in Figure 4. We employed a TurtleBot 2 robotic platform [46] specifically adapted for the ObjectNav problem. This adaptation replicates the agent’s characteristics during training. Among the modifications, we added a mast to the TurtleBot 2, raising the camera to 1.25 m to match the simulation setup. The camera used is an Orbbec Astra with depth perception. Since our SemNav model requires a semantic segmentation sensor, for real-world testing, we integrated the ESANet model [47], pre-trained on the NYUv2 dataset [42]. This model takes the robot’s egocentric vision as input in RGB+Depth format and generates a semantic segmentation image, which is then fed into our SemNav model.
The embedding of our SemNav model and the state-of-the-art PirlNav [48] in the robotic platform was done using the ROS4VSN library [24], which allows for the integration of VSN solutions into real robots using ROS. We conducted navigation experiments toward five specific object classes: chair, bed, toilet, television monitor, and sofa. For each episode, an object category was selected as the navigation target, and the navigation started from a predefined location and orientation. These starting points were chosen to ensure that: 1) there was no direct visibility of the target object; and 2) there was a minimum distance of 10 meters to the object. The evaluation metric used was the SR, where a navigation episode is considered successful if the robot infers the stop action while being within one meter of the target object. A failure is recorded if the robot collides and cannot proceed or if it requires more than 225 actions to complete the task.
@@
Figure 4: Top-down view of the house where ObjectNav experiments were conducted for five object categories.
We release all code for training and evaluating our models, both in simulated environments and in the real world, in the following anonymous repository.
@@ IV-B1 Ablation study
In this analysis, we aim to present the impact of the following aspects on the performance of our SemNav model: 1) We test a SemNav configuration that uses only semantic segmentation as visual input—SemNav-Only Semantic Segmentation (SemNav-OS); 2) We add a standard RGB visual input to the SemNav model, which passes through its corresponding visual encoder (a ResNet 50) to be concatenated with the rest of input features—SemNav-RGB + Semantic Segmentation (SemNav-RGBS); 3) Each of these configurations is tested with semantic segmentation using 1630 and 40 categories, as provided in the SemNav dataset; and 4) For each test, we train the model either from randomly initialized weights or by pre-initializing them using DINO [44], since this initialization strategy has demonstrated a significant impact on previous VSN models, e.g., PirlNav [48]. In total, we propose an evaluation of up to eight different configurations for our SemNav model.
@@
SemNav 1630 OS 65.1 0.29 57.11 2.99
OS+DINO 68.4 0.30 50.39 2.75
RGBS 69.2 0.31 50.59 2.73
RGBS+DINO 70 0.32 39.17 2.69
SemNav 40 OS 72.9 0.34 43.25 2.55
OS+DINO 73 0.34 48.49 2.61
RGBS 74.3 0.35 40.70 2.50
RGBS+DINO 76.2 0.36 38.50 2.36
@@
TABLE I: Evaluation of the performance of different configurations of our SemNav model. We report the metrics: Succes Rate (SR), Shortest Path Length (SPL), average number of collisions (C), and average distance to the goal (DTG).
@@
Table I presents the performance of all these configurations on the SemNav-dataset validation set. In light of the results obtained in this study, our model SemNav +RGBS has provided the best performance. It is interesting to highlight that all models trained with 40-category semantic segmentations have outperformed those trained with 1630-category semantic segmentations. We believe this is primarily due to the annotation issues present in the 1630-category set, as detailed in Section III-A, and the fact that the semantic information necessary for successful navigation is already included in the 40 categories, making the 1630-category segmentation unnecessarily complex. Lastly, our SemNav model does not appear to be significantly affected by initialization with DINO [44], unlike other state-of-the-art models. In our case, DINO initialization slightly improves the results.
@@
This ablation study also demonstrates that semantic segmentation is the key factor impacting model performance. Comparing any of our models with and without RGB input—i.e., OS vs. RGBS—further confirms this observation. For instance, in terms of SR, for the SemNav 40 dataset, the SemNav model improves from 72.9 to 74.3 when using the OS and RGBS versions, respectively. We conclude that these results validate the suitability of the SemNav approach for the VSN problem. Adding RGB information to the model consistently enhances its performance, and for real-world navigation, this would be the optimal architecture to embed in a robotic platform, as we will demonstrate in Section IV-C.
IV-B2 Comparison with the state of the art

### v2

@@ III-B SemNav model
The architecture depicted in Figure 3 comprises several key components. To process the semantic segmentation input, we first transform it into an RGB format. Each pixel is assigned a semantic segmentation label, which is then mapped to a specific color in the RGB space using the semantic segmentation sensor integrated into the SemNav dataset. The encoded input is subsequently processed by a ResNet-50 architecture [17], which extracts high-level visual features for downstream tasks. This process is formalized in the following equation:
$h^{res}_{i}=f_{\text{RESNET}}(o^{\text{semantic}}_{i};\theta_{\text{RESNET}})\in\mathbb{R}^{d_{\text{res}}},$ (2)
where $h^{\text{res}}_{i}$ represents the extracted visual feature vector at position $p_{i}$, $f_{\text{RESNET}}(\cdot)$ denotes the ResNet-50 feature extraction function, $o^{\text{semantic}}_{i}$ is the semantic segmentation input at position $p_{i}$, and $\theta_{\text{RESNET}}$ encodes the learnable parameters of the ResNet-50. Here, the output dimensionality of the ResNet-50 is $d_{\text{res}}=512$, meaning that the resulting feature vector belongs to a 512-dimensional real-valued space.
For the experimental evaluation, we also used architectures that incorporate multiple visual modalities. Technically, we are able to integrate both semantic segmentation and RGB images as inputs. In this case two parallel ResNet-50 backbones are employed, each dedicated to processing one modality independently. This architectural design was selected not only for its favorable balance between computational efficiency and performance, but also to ensure a fair and consistent comparison with other ObjectNav models that adopt the same configuration, thereby enabling a more equitable evaluation framework.
In our experiments, we evaluate two types of initialization for the ResNet50: (1) a random initialization and (2) an initialization using DINO [6].
The input corresponding to the target object category is represented by its class index $c_{i}$ and processed through an embedding layer, as shown in the following equation:
$h^{\text{goal}}_{i}=W_{\text{goal}}[c_{i}]\in\mathbb{R}^{d_{\text{goal}}},$ (3)
where $h^{\text{goal}}_{i}$ denotes the embedded representation of the target category at timestep $i$, and $W_{\text{goal}}\in\mathbb{R}^{m\times d}$ represents the learnable embedding matrix, where $m=6$ is the number of object categories and $d_{\text{goal}}=32$ is the embedding dimension. The bracket notation $\cdot$] indicates the selection of the corresponding row from the embedding matrix associated with category $c_{i}$.
The position and orientation observations, both expressed relative to the starting position, are processed through independent fully connected layers. The position observation is projected from $\mathbb{R}^{2}$ to a $d_{\text{pos}}$-dimensional embedding space, and the orientation observation, represented by the cosine and sine of the relative heading angle, is projected from $\mathbb{R}^{2}$ to a $d_{\text{ori}}$-dimensional embedding space. This process is formalized in the following equations:
$h^{pos}_{i}=W_{pos}\,o^{pos}_{i}+b_{pos}\in\mathbb{R}^{d_{\text{pos}}},\quad h^{ori}_{i}=W_{ori}\,o^{ori}_{i}+b_{ori}\in\mathbb{R}^{d_{\text{ori}}}\;.$ (4)
Here, $o^{pos}_{i}\in\mathbb{R}^{2}$ and $o^{ori}_{i}\in\mathbb{R}^{2}$ denote the position and orientation observations at timestep $i$, respectively. The learnable parameters $W_{pos},W_{ori}\in\mathbb{R}^{d_{\text{pos}}\times 2},\mathbb{R}^{d_{\text{ori}}\times 2}$ and $b_{pos},b_{ori}\in\mathbb{R}^{d_{\text{pos}}},\mathbb{R}^{d_{\text{ori}}}$ correspond to the weights and biases of the linear embedding layers. In our implementation, $d_{\text{pos}}=d_{\text{ori}}=32$, and the resulting feature vectors $h^{pos}_{i},h^{ori}_{i}$ are then concatenated with the other observation embeddings before being passed to the recurrent module.
To additionaly incorporate the agents previous action into the policy input, an embedding of the action taken at the previous timestep is included. Let $a_{i-1}$ denote the previous action and $W_{act}\mathbb{R}^{n_{act}\times d_{act}}$ be the learnable embedding matrix where $n_{act}=6$ is the number of discrete actions and $d_{act}$ is the action embedding dimensionality. The previous action embedding is obtained as:
$h^{\text{act}}_{i}=W{\text{act}}[a_{i-1}]\in\mathbb{R}^{d_{\text{act}}}.$ (5)
All the processed observations, including the visual embedding from the ResNet-50, the object goal embedding, and the embeddings from the position and orientation, are concatenated into a single feature vector at each timestep:
$x_{i}=\big[h^{res}_{i}\,;\,h^{emb}_{i}\,;\,h^{pos}_{i}\,;\,h^{ori}_{i}\,;\,h^{act}_{i}\big]\in\mathbb{R}^{d_{x}},$ (6)
where $d_{x}=d_{res}+d_{emb}+d_{pos}+d_{ori}+d_{act}$ denotes the total dimensionality of the concatenated input, and [;][\,;\,] represents vector concatenation.
This vector $x_{i}$ is then passed through a Gated Recurrent Unit (GRU) [9] to capture temporal dependencies and provide the model with a memory of past navigation decisions. In our implementation, the GRU has two recurrent layers ($\text{num\_recurrent\_layers}=2$) and a hidden size of $d_{h}=2048$, as specified in the configuration:
$h_{\text{i}}=\text{GRU}(x_{i},h_{\text{i}-1};\theta_{\text{i}})\in\mathbb{R}^{d_{h}},$ (7)
where $h_{i}\in\mathbb{R}^{d_{h}}$ is the hidden state of the GRU at timestep $i$, $d_{h}=2048$ denotes the dimensionality of the hidden layer, and $\theta_{\text{GRU}}$ represents the learnable parameters of the GRU across both layers. The GRU allows the model to condition its current action predictions on past observations without the need for an explicit map, enabling autonomous, map-free navigation. This way, the recurrent hidden state $h_{i}$ conditions the model’s action predictions, by taking into account the history of past observations and navigation actions. The GRU output is subsequently used to predict the discrete action distribution that forms the output of our policy.
The hidden state $h_{i}\in\mathbb{R}^{d_{h}}$ produced by the GRU at timestep $i$ is subsequently fed into a multi-layer perceptron (MLP) to produce the unnormalized action logits. Let $n_{a}$ denote the number of discrete actions; in our case, $n_{a}=6$, because our action space is $\mathcal{A}=\{\textsc{turn\_left},\textsc{turn\_right},\textsc{move\_forward},\allowbreak\textsc{move\_backward},\textsc{stop}\}$. The process can be formalized as follows:
$z_{i}=f_{\text{MLP}}(h_{i};\theta_{\text{MLP}})\in\mathbb{R}^{n_{a}},$ (8)
where $f_{\text{MLP}}(\cdot)$ represents the multi-layer perceptron with learnable parameters $\theta_{\text{MLP}}$, and $z_{i}$ are the logits corresponding to each discrete action.
These logits are then passed through a categorical distribution to obtain the action probabilities:
$\pi(a_{p_{i}}\mid o_{p_{i}})=\text{Categorical}(\text{softmax}(z_{i}))\in\mathbb{R}^{n_{a}},$ (9)
where $\pi(a_{p_{i}}\mid o_{p_{i}})$ defines the probability distribution over the $n_{a}$ possible actions given all observations up to timestep $i$, and the softmax function ensures a valid probability distribution.
In summary, the recurrent hidden state $h_{i}$ encodes the temporal context and is transformed by the MLP to produce logits, which are subsequently normalized by the softmax function to define a categorical policy over the discrete action space.
The complete training procedure for the SemNav visual navigation policy is summarized in Algorithm 1, which formalizes the imitation learning process described above. The algorithm details the iterative optimization of the network parameters $\theta$ using expert demonstrations from the dataset $\mathcal{T}$, encompassing the visual encoding, feature embedding, recurrent processing, and action prediction steps at each timestep. This end-to-end learning framework enables the agent to acquire robust navigation policies that leverage semantic segmentation information to navigate efficiently toward target object categories in previously unseen environments.
Algorithm 1 Imitation Learning for SemNav Visual Navigation Policy
[…]
1: Initialize $\theta$ randomly or with DINO pre-training for $\theta_{\text{RESNET}}$
2:for epoch $e=1$ to $E$ do
3:  Shuffle trajectories $\mathcal{T}$
4:  for each mini-batch of $B$ trajectories do
5:   Initialize gradient accumulator $\nabla_{\theta}\mathcal{L}\leftarrow 0$
6:   for all trajectory $\tau=(o_{p_{0}},a_{p_{0}},\dots,o_{p_{T}},a_{p_{T}})\in\text{batch}$ do
7:    Initialize recurrent state $h_{0}\in\mathbb{R}^{d_{h}}$
8:    for all timestep $t=0$ to $T$ do
9:     // Visual encoding
10:     $h^{\text{res}}_{t}\leftarrow f_{\text{RESNET}}(o^{\text{semantic}}_{t};\theta_{\text{RESNET}})$
11:     // Goal, position, previous action and orientation embeddings
12:     $h^{\text{goal}}_{t}\leftarrow W_{\text{goal}}[c_{i}]$
13:     $h^{\text{pos}}_{t}\leftarrow W_{\text{pos}}\,o^{\text{pos}}_{t}+b_{\text{pos}}$
14:     $h^{\text{ori}}_{t}\leftarrow W_{\text{ori}}\,o^{\text{ori}}_{t}+b_{\text{ori}}$
15:     $h^{\text{act}}_{t}\leftarrow W{\text{act}}[a_{i-1}]$
16:     // Concatenate features
17:     $x_{t}\leftarrow[h^{\text{res}}_{t};h^{\text{goal}}_{t};h^{\text{pos}}_{t};h^{\text{ori}}_{t}]$
18:     // Recurrent processing
19:     $h_{t}\leftarrow\text{GRU}(x_{t},h_{t-1};\theta_{\text{GRU}})$
20:     // Action prediction
21:     $z_{t}\leftarrow f_{\text{MLP}}(h_{t};\theta_{\text{MLP}})$
22:     $\pi_{\theta}(a_{p_{t}}\mid o_{p_{t}})\leftarrow\text{Categorical}(\text{softmax}(z_{t}))$
23:     // Accumulate loss
24:     $\nabla_{\theta}\mathcal{L}\leftarrow\nabla_{\theta}\mathcal{L}-\nabla_{\theta}\log\pi_{\theta}(a_{p_{t}}\mid o_{p_{t}})$
25:    end for
26:   end for
27:   // Update parameters
28:   $\theta\leftarrow\theta-\alpha\cdot\nabla_{\theta}\mathcal{L}/B$
29:  end for
30:end for
31:return Trained policy parameters $\theta$
@@ IV-A Experimental settings and evaluation metrics
Our main goal in this work is to develop VSN solutions capable of navigating efficiently both in the virtual environments and in the real world. For this purpose, we have designed two types of experiments.
@@
For this experimental evaluation, we addressed the ObjectNav task detailed in the Habitat 2023 challenge [42]. We used the official HM3D [45] training and validation sets. For our SemNav model we employed the two semantic segmentation sensors detailed in Section III-A. For evaluation metrics, we followed the standard proposals for the ObjectNav problem [42]. Specifically, these metrics include: 1) the Success Rate (SR), defined as the percentage of navigation episodes successfully completed by the model; 2) the Shortest Path Length (SPL), which compares the distance traveled by the agent during the episode with the shortest possible path length to the target object; 3) the average number of collisions the agents experience with the environment (C); and 4) the average distance to the goal (DTG), which is the mean distance of the agent from the nearest target object at the end of a navigation episode. A navigation episode is considered successful when the agent samples the stop action within 1 meter of the target object it is tasked to find.
@@
To assess the performance of our models in the real world, experiments were conducted in three different houses, whose floor plans are shown in Figure 4. That is, the SemNav model, trained entirely in simulation, is now evaluated in the real world—within three unseen environments that present a significant domain gap. This type of experiment is essential to assess whether our proposed SemNav approach can generalize effectively and navigate reliably in real-world conditions.
We employed a TurtleBot 2 robotic platform [23] specifically adapted for the ObjectNav problem. This adaptation replicates the agent’s characteristics during training. Among the modifications, we added a mast to the TurtleBot 2, raising the camera to 1.25 m to match the simulation setup. The camera used is an Orbbec Astra with depth perception. Since our SemNav model requires a semantic segmentation sensor, for real-world testing, we integrated the ESANet model [33], pre-trained on the NYUv2 dataset [35]. This model takes the robot’s egocentric vision as input in RGB+Depth format and generates a semantic segmentation image, which is then fed into our SemNav model. Each segmentation inference incurs an additional overhead of approximately 11ms per frame when running on an NVIDIA RTX4080. In these real-world experiments, we start with a detailed comparative study of performance between our proposed SemNav model and state-of-the-art VSN approaches, incorporating the PirlNav model [26] into the evaluation. This comparative experiment was carried out in the house 1, whose floor plan can be seen in Figure 4a. We analyze in house 1, the main types of errores that our models exhibit, as well as for the state-of-the-art PirlNav approach. Then, we extend the whole experimental comparison to houses 2 and 3, see Figures 4b and 4c, respectively. Our best-performing model in simulation, SemNav-RGBS+DINO $\rightarrow$ RL, was evaluated in all these houses.
The embedding of our SemNav model and the state-of-the-art PirlNav [26] in the robotic platform was done using the ROS4VSN library [16], which allows for the integration of VSN solutions into real robots using ROS. We conducted navigation experiments toward five specific object classes: chair, bed, toilet, television monitor, and sofa. For each episode, an object category was selected as the navigation target, and the navigation started from a predefined location and orientation. These starting points were chosen to ensure that: 1) there was no direct visibility of the target object; and 2) there was a minimum distance of 10 meters to the object. The evaluation metric used was the SR, where a navigation episode is considered successful if the robot infers the stop action while being within one meter of the target object. A failure is recorded if the robot collides and cannot proceed or if it requires more than 225 actions to complete the task. In the real world, $N_{\text{max}}=225$, as real environments typically offer less navigable surface compared to simulation.
@@
(a) House 1
(b) House 2
(c) House 3
Figure 4: Top-down views of the houses where ObjectNav real-world experiments were conducted for five object categories.
We release all code for training and evaluating our models, both in simulated environments and in the real world, in the following repository.
@@ IV-B1 Ablation study
In this analysis, we aim to present the impact of the following aspects on the performance of our SemNav model: 1) We test a SemNav configuration that uses only semantic segmentation as visual input—SemNav-Only Semantic Segmentation (SemNav-OS); 2) We add a standard RGB visual input to the SemNav model, which passes through its corresponding visual encoder (a ResNet 50) to be concatenated with the rest of input features—SemNav-RGB + Semantic Segmentation (SemNav-RGBS); 3) Each of these configurations is tested with semantic segmentation using 1630 and 40 categories, as provided in the two SemNav dataset sensors; and 4) For each test, we train the model either from randomly initialized weights or by pre-initializing them using DINO [6], since this initialization strategy has demonstrated a significant impact on previous VSN models, e.g., PirlNav [26]. In total, we propose an evaluation of up to eight different configurations for our SemNav model.
@@
SemNav 1630 PirlNav (RGB) 60.9 0.26 45.74 3.17
SemNav-OS 65.1 0.29 57.11 2.99
SemNav-OS+DINO 68.4 0.30 50.39 2.75
SemNav-RGBS 69.2 0.31 50.59 2.73
SemNav-RGBS+DINO 70 0.32 39.17 2.69
SemNav 40 PirlNav (RGB) 60.9 0.26 45.74 3.17
SemNav-OS 72.9 0.34 43.25 2.55
SemNav-OS+DINO 73 0.34 48.49 2.61
SemNav-RGBS 74.3 0.35 40.70 2.50
SemNav-RGBS+DINO 76.2 0.36 38.50 2.36
@@
TABLE I: Evaluation of the performance of different configurations of our SemNav model. We report the metrics: SR, SPL, C, and DTG.
@@
Table I presents the performance of all these configurations on the SemNav-dataset validation set. In light of the results obtained in this study, our model SemNav-RGBS has provided the best performance. It is interesting to highlight that all models trained with 40-category semantic segmentations have outperformed those trained with 1630-category semantic segmentations. We believe this is primarily due to the annotation issues present in the 1630-category set, as detailed in Section III-A, and the fact that the semantic information necessary for successful navigation is already included in the 40 categories, making the 1630-category segmentation unnecessarily complex. Lastly, our SemNav model does not appear to be significantly affected by initialization with DINO [6], unlike other state-of-the-art models. In our case, DINO initialization slightly improves the results.
@@
This ablation study also demonstrates that semantic segmentation is the key factor impacting model performance. Comparing any of our models with and without RGB input—i.e., SemNav-OS vs. SemNav-RGBS—further confirms this observation. For instance, in terms of SR, for the SemNav 40 dataset, the SemNav model improves from 72.9 to 74.3 when using the SemNav-OS and SemNav-RGBS versions, respectively. We conclude that these results validate the suitability of the SemNav approach for the VSN problem. Adding RGB information to the model consistently enhances its performance, and for real-world navigation, this would be the optimal architecture to embed in a robotic platform, as we will demonstrate in Section IV-C.
IV-B2 Semantic segmentation precision impact
The training of the SemNav models has been conducted using the semantic segmentation information provided by our novel semantic sensors in the SemNav dataset. These semantic segmentations are derived from manually annotated data, which provides high accuracy. However, we consider it important to evaluate the impact that the quality of semantic segmentation may have on navigation performance. This section presents a study on the impact of using semantic segmentations of varying quality through the following experiments. Technically, we propose to tweak the original semantic segmentation quality in the SemNav sensor 40 by injecting three types of noise to mimic imperfections that could appear in a real-world application:
  1. 1.
    Category misclassification noise: This noise model introduces semantic confusion by injecting probabilistic label errors. Each semantic class is assigned a probability of being misclassified as another class. For example, pixels belonging to the category chair may be incorrectly labeled as table or floor, emulating the typical misclassification patterns observed in automatic segmentation systems when differentiating visually similar objects.
  2. 2.
    Boundary noise: This perturbation degrades boundary accuracy by introducing spatial and temporal distortions along segmentation edges. Unlike the sharp, well-defined contours in ground-truth annotations, this noise generates irregular, jagged boundaries that fluctuate across consecutive frames, effectively modeling the uncertainty and temporal inconsistency observed in real-world semantic segmentation systems, particularly near object transitions and occlusion regions.
  3. 3.
    Spatially localized perturbations: This noise introduces discrete, spatially confined artifacts by randomly sampling patches of incorrect labels across the image. These perturbations emulate localized segmentation failures, such as misclassified regions, spurious detections, or missing object parts. These are typical artifacts in real-world semantic segmentation systems caused by challenging illumination, texture ambiguities, or sensor-induced noise.
In this study on the influence of semantic segmentation accuracy, we focus primarily on category misclassification noise, as it represents the most prevalent error type in both simulated (arising from dataset misannotations) and real-world semantic segmentation systems. We prioritize this noise type in our analysis before extending the evaluation to boundary and spatially localized perturbations.
SR with sensor errors 0% 1% 10% 30%
SR 76.2 74.5 73.3 69.9
SPL 0.36 0.353 0.337 0.287
TABLE II: Evaluation of SR and SPL under different levels of Category misclassification sensor noise.
Table II shows the impact that this type of noise has on the metrics used to evaluate navigation performance. Technically, in this experiment we have used noisy semantic segmentation sensors with varying misclassification probabilities of $0\%$, $1\%$, $10\%$, and $30\%$. The noise injection process operates as follows: when a batch of images is received, each semantic category within the batch has a certain probability of being incorrectly relabeled as a different category from the 40 available classes. Critically, when a category is selected for misclassification, all pixels belonging to that category are simultaneously relabeled to the same alternative category. This probabilistic misclassification is applied at the batch level, meaning the category substitutions vary across different batches and do not persist throughout the entire navigation episode. Consequently, the spatial structure and boundaries of the segmentation remain intact, as they are directly derived from the ground truth annotations; only the semantic labels are altered coherently across all pixels of each affected category. Figure 5c illustrates the visual effect of category misclassification noise at a $10\%$ error rate. As shown, the wall pixels have been incorrectly relabeled, resulting in a different color assignment compared to the ground-truth semantic segmentation depicted in Figure 5b.
Figure 5 (c) illustrates the visual effect of this category misclassification noise.
For the specific evaluation of the impact of the noisy semantic segmentations, the performance was assessed using SR, SPL, C, and DTG metrics with our SemNav-RGBS+DINO model trained via IL. These metrics are reported using 2000 navigation trajectories. The results of this experiment are presented in Table II. As the misclassification rate of the semantic segmentation sensor increases, both SR and SPL metrics degrade, highlighting the sensitivity of navigation performance to segmentation quality. This experiment yields two main conclusions. First, it underscores the critical role of semantic information in navigation tasks: higher semantic segmentation accuracy directly correlates with improved navigation performance. Second, when deploying models such as SemNav in real-world scenarios—where accurate or manually annotated semantic segmentations are unavailable—a degradation in navigation performance might be expected.
What happens when the remaining types of noise are taken into consideration? In this second experiment, each noise type was independently injected into the semantic segmentation sensor, as well as a combined configuration including all noise types. Results over 2000 evaluation trajectories are summarized in Table III.
Noise Type SR (%) SPL C DTG
No Noise 76.2 0.36 38.50 2.36
Boundary noise 74.5 0.352 38.6 2.47
10% Category misclassification 73.3 0.337 40.73 2.5
Spatially localized perturbations 68.95 0.288 44.2 2.79
All Noises Combined 64.05 0.252 48.49 3.07
TABLE III: Evaluation of SR, SPL, C, and DTG under different types of semantic segmentation sensor noise.
As shown in Table III, introducing noise into the semantic segmentation negatively impacts the navigation performance of the SemNav model, regardless of the noise type. Boundary segmentation errors cause a moderate decrease in performance, whereas category misclassification noise and spatially localized perturbations have a more pronounced effect.
Boundary segmentation errors cause a moderate decrease in performance, whereas category misclassification noise and spatially localized perturbations have a more pronounced effect. This is because abrupt changes in object geometry do not fundamentally alter the semantic structure of the navigated space, whereas category misclassifications or spurious category labels can mislead the agent into reasoning that it is in a semantically different location. For instance, if the agent observes pixels or objects labeled as sofa within a bathroom, this semantic inconsistency significantly disrupts the navigation process.
When all noise types are combined, the model’s performance degrades substantially, highlighting the importance of accurate semantic segmentation for optimal performance in visual navigation tasks. The qualitative impact of the proposed noise models is illustrated in Figure 5, which shows examples of semantic segmentations under different noise configurations. The applied perturbations generate simulated artifacts that could resemble those produced by automatic semantic segmentation models in real-world conditions. This similarity is particularly evident in the comparison between images (f) and (g), where the noise introduced by the actual segmenter is contrasted with the combined effect of the three proposed noise types.
The segmentation in (f) produced by ESANet [33] exhibits category misclassification noise, observable in the floor, ceiling, and wall paintings; boundary errors, evident in the less sharp floor edges compared to noise-free simulation; and spatially localized perturbations, manifested as spurious object categorizations within other objects, such as on the sofa or paintings. Image (g) artificially replicates this behavior by combining all noise types, showing misclassifications (e.g., ceiling), boundary degradation across scene objects, and spatially localized perturbations in the central region.
Figure 5: Qualitative results in the simulation environment for the different noisy semantic segmentation sensors. (a) RGB sensor output overlaid on the ground truth, (b) semantic segmentation sensor with no noise, (c) semantic segmentation missclassification noise with 10% error, (d) spatially localized perturbations noise error, (e) Border Error noise, (f) ESANet output, (g) semantic segmentation missclassification noise, spatially localized perturbation noise error and border error noise combined.
IV-B3 Resources analysis
Our SemNav model needs to perform a semantic segmentation within the inference loop, therefore an additional temporal and computational overhead is introduced during inference. In this experiment we perform a detailed runtime analysis, to show the impact of the use of a semantic segmentation module in our pipeline. Technically, as in the real-world experiments, we integrate in SemNav the ESANet [33] model. This semantic segmentation model is utilized solely to generate the semantic segmentation image that serves as input for the SemNav-OS and SemNav-RGBS pipelines. Consequently, the total time per step and memory requirements must account for both the latency of the segmentation model and the latency of the action decision model. Table IV presents a temporal breakdown per inference step, detailing the inference times for both the action models and the segmentation model. Table V summarizes the GPU memory requirements for each component (values are representative measurements obtained on an NVIDIA RTX4080).
A robotic navigation model intended for real-world deployment must exhibit low-latency temporal response, ensuring the user perceives an active system without significant delays in action execution. For this reason, the temporal analysis presented in Table IV reports our model’s response times. Since the system must infer a semantic segmentation image from an RGB input, some computational overhead is introduced, which could limit practical applicability. However, as shown, although this overhead is considerably larger than the time required by the model to decide the action once all inputs are processed, the overall response times remain sufficiently low to enable more than 110 inferences per second. In practice, this inference time is negligible compared to the time required for the robot to physically execute each action, which is significantly longer than the model’s inference time.
TABLE IV: Temporal breakdown per inference step (representative values, ms). SemNav-OS and SemNav-RGBS are action decision models; ESANet [33] is the semantic segmentation model. “Total” = action model + semantic segmentation model.
Model Model Inf. (ms) Semantic Segmentation Inf. (ms) Total per step (ms) Approx. FPS
SemNav-OS 1 7.5 8.5 117
SemNav-RGBS 1.5 7.5 9 111
Furthermore, a memory-efficient navigation model is more suitable for deployment across diverse environments and platforms. Table V presents the GPU memory footprint of each model evaluated in this article. The reported sizes demonstrate that these models maintain a compact memory profile, making them compatible with a wide range of graphics hardware.
TABLE V: Model size in GPU memory (MB). The reported sizes correspond to the model parameters loaded on GPU.
Model Model size on GPU (MB) Notes
SemNav-OS 188.54 ResNet-50 backbone + head
SemNav-RGBS 228.79 Dual backbones (RGB + seg.)
ESANet  [33] 181.27 Semantic segmentation module
Power consumption is also a limiting factor for model deployment, particularly when integration into compact, low-power devices is desired. Table VI reports the power consumption of navigation models on an NVIDIA RTX4080 graphics card. As expected, training consumes significantly more power than inference, while inference power requirements remain modest and manageable for practical deployment.
TABLE VI: Power consumption during training and inference (representative values, W). Measurements obtained on an NVIDIA RTX4080.
Model Training (W) Inference (W)
SemNav-OS 115 13
SemNav-RGBS 140 30
Finally, Table VII presents the hyperparameters used for training the SemNav-RGBS model via IL. These parameters were employed on 8 NVIDIA A100 GPUs over 10 days using the SemNav 40 dataset. These specifications highlight the computational cost of training and facilitate the reproducibility of our results.
TABLE VII: Training configuration and hyperparameters used for Behavior Cloning.
Parameter Value
Number of GPUs 8
Number of environments per GPU 16
Rollout length 64
Number of mini-batches per epoch 2
Optimizer Adam
Learning rate scheduler Cyclic LR (exp_range)
Base learning rate $1\times 10^{-5}$
Maximum learning rate $1\times 10^{-3}$
Step size up 2000
Exponential decay factor ($\gamma$) 0.99994
DDPIL sync fraction 0.6
IV-B4 Comparison with the state of the art